OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos(5π/12) I know someone knows how to do this! if you could explain too, that'd be great!

OpenStudy (jdoe0001):

\(\bf cos\left(\frac{5\pi}{12}\right)\\ \quad \\ \quad \\ -------------\\ \cfrac{5}{12}\times 2\implies \cfrac{5}{6}\\ \quad \\ cos\left(\frac{5\pi}{12}\right)\implies cos\left(\cfrac{\frac{5\pi}{6}}{2}\right)\\ \quad \\ -------------\\ cos\left(\frac{\theta}{2}\right) = \pm \sqrt{\cfrac{1+cos(\theta)}{2}}\\ \quad \\ cos\left(\cfrac{\frac{5\pi}{6}}{2}\right) = \pm \sqrt{\cfrac{1+cos\left(\frac{5\pi}{6}\right)}{2}} \) now, check your Unit Circle to see what's the \(\bf cos\left(\frac{5\pi}{6}\right)\)